# Constraining interacting dark energy models with black hole superradiance

### v1

@@ 3.3 Statistical framework
To formalize this exclusion principle in a Bayesian context, we follow the statistical framework developed by Hoof et al. (2020) [Hoof:2024quk] but adapt it to our specific model and parameter space. The central idea is to compute the posterior probability for our model parameters $\boldsymbol{\alpha}$ given the observational data $D$ from a BH measurement.
@@
$p(\boldsymbol{\alpha}|D)$ $=p(\boldsymbol{\alpha})\int\!p(\boldsymbol{\beta}_{\mathrm{BH}}|D)\,p(\tilde{a}|M,\boldsymbol{\alpha})\,\mathrm{d}\boldsymbol{\beta}_{\mathrm{BH}}\,.$ (3.12)
@@
In practice, observational analyses provide a set of $N$ equally-weighted posterior samples, $\{\boldsymbol{\beta}_{\mathrm{BH}}^{i}=(M^{i},\tilde{a}^{i})\}_{i=1}^{N}$, for the BH parameters. The integral in Eq. (3.12) can therefore be accurately approximated via Monte Carlo integration as
@@
$p(\boldsymbol{\alpha}|D)\approx\frac{p(\boldsymbol{\alpha})}{N}\sum_{i=1}^{N}\Theta(\tilde{a}^{\text{crit}}(M^{i};\boldsymbol{\alpha})-\tilde{a}^{i})\,.$ (3.13)
@@
The sum effectively counts the fraction of posterior samples for the BH that are consistent with (i.e., not excluded by) the given set of model parameters $\boldsymbol{\alpha}$ [Hoof:2024quk]. Since the superradiance rate $\Gamma_{\mathrm{SR}}$, and thus $\tilde{a}^{\text{crit}}$, depends on the properties of the ultralight boson derived from the specific IDE model (such as its effective mass, which is a function of $\boldsymbol{\alpha}$), this statistical framework allows us to translate observational data of highly spinning BHs into exclusion regions in the parameter space $\boldsymbol{\alpha}$ of the IDE model under consideration.
4 Model I: DE as a dark fifth force mediator of the DM sector
@@
We now introduce the first of our specific field-theoretic models, building upon the general IDE framework discussed in Section 2. In this scenario, the DE scalar field acts as a dark fifth force mediator of the DM sector, assuming both are ultralight scalar fields. This interaction alters the DM effective mass based on the background DE field, thereby modifying the superradiance potential of the DM particles themselves around spinning BHs and enabling constraints on the DE-DM coupling.
4.1 Trilinear coupled quintessence model
@@
We begin by considering a coupled quintessence model where a DE scalar field, $\phi$, is coupled to an ultralight DM field  [^2] The assumption that DM is an ultralight scalar is motivated by the naturalness of the model [Archidiacono:2022iuu]. Theoretical consistency requires that the parameters of the $\phi$-potential are at least as large as the radiative corrections induced by the interaction with DM. This leads to the condition $m_{\phi}^{2}\gtrsim\frac{\beta}{(4\pi)^{2}}\frac{m_{\chi}^{4}}{M_{\mathrm{Pl}}^{2}}\,.$ (4.1) Imposing $m_{\phi}\lesssim H_{0}$ yields an upper bound on the DM mass, $m_{\chi}\lesssim\beta^{-1/4}\left(4\pi m_{\varphi}M_{\mathrm{Pl}}\right)^{1/2}\simeq 0.02\ \mathrm{eV}\left(\frac{0.01}{\beta}\right)^{1/4}\left(\frac{m_{\varphi}}{H_{0}}\right)^{1/2}\,.$ (4.2) Therefore, a relativistic field-theoretical description of the dark fifth force mediator without fine-tuning requires the DM to be an ultralight boson. In this mass range, the DM candidate could be an axion or axion-like particle produced via the misalignment mechanism or other non-thermal processes., $\chi$, via a trilinear interaction term [Farrar:2003uw, Archidiacono:2022iuu, Bottaro:2023wkd, Bottaro:2024pcb]. The Lagrangian for this model is given by
$\mathcal{L}=-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m_{\phi}^{2}\phi^{2}-\kappa\phi\chi^{2}\,$, (4.3)
@@
$\mathcal{L}=-\frac{1}{2}(\partial\chi)^{2}-\frac{1}{2}m_{\chi}^{2}(s)\chi^{2}-\frac{1}{2G_{s}}\left[(\partial s)^{2}+m_{\phi}^{2}s^{2}\right]\,.$ (4.5)
@@
Models of this type, where an ultralight scalar field (which may or may not be the DE field) mediates a long-range fifth force exclusively within the DM sector, have been extensively studied in the cosmological literature [Wetterich:1994bg, Amendola:1999er, Farrar:2003uw, Archidiacono:2022iuu, Bottaro:2023wkd, Bottaro:2024pcb]. These studies typically focus on constraining the model parameters using cosmological observations. Indeed, existing cosmological data, primarily from the Cosmic Microwave Background [Archidiacono:2022iuu] and large-scale structure surveys [Bottaro:2023wkd, Bottaro:2024pcb], already place stringent constraints on the dimensionless coupling strength, typically finding $\beta\lesssim\mathcal{O}(0.01)$. Our work complements these cosmological probes by exploring constraints from the distinct physical regime of BH superradiance.
4.2 Effective mass correction
The coupling between the DE field $\phi$ and the DM field $\chi$ results in the mass of the DM particle becoming dependent on the local background value of the DE field, $\bar{s}$. This gives rise to an effective mass for DM, which, for our model, is given by
$m_{\chi}^{2}(\bar{s})=m_{\chi}^{2}(1+2\bar{s}),$ (4.8)
where $m_{\chi}$ is the bare mass of the DM particle.
@@ 4.2 Effective mass correction
In principle, the value of the background field $\bar{s}$ evolves with cosmic time, governed by its equation of motion [Archidiacono:2022iuu],
$\ddot{\bar{s}}+3H\dot{\bar{s}}+\left(m_{\phi}^{2}\bar{s}+4\pi G_{N}\beta\frac{\partial\log m_{\chi}(\bar{s})}{\partial\bar{s}}\bar{\rho}_{\chi}\right)=0\,.$ (4.9)
A full analysis would require solving this equation over cosmic history, assuming a fixed initial value (for a scalar mass $m_{\phi}\sim H_{0}$, one finds $\bar{s}_{\mathrm{ini}}\sim\mathcal{O}(1)$). However, since the BH sample we selected formed very recently—roughly $\mathcal{O}(10-10^{2})\,\mathrm{Myr}$ ago, and the superradiance time scale is too short for significant change of $\bar{s}$, we can approximate $\bar{s}$ with its constant, present-day value $\bar{s}_{0}$ when applying superradiance constraints.
@@
$V(\bar{s}_{0})\approx\rho_{\mathrm{DE}}\approx 0.7\times 3H_{0}^{2}M_{\mathrm{Pl}}^{2}\,$, (4.10)
@@
$V(\bar{s})=\frac{1}{2}m_{\phi}^{2}\bar{\phi}^{2}=\frac{M_{\mathrm{Pl}}^{2}m_{\phi}^{2}\bar{s}^{2}}{\beta}.$ (4.11)
@@
$\bar{s}_{0}\approx 1.45\frac{H_{0}}{m_{\phi}}\beta^{1/2}.$ (4.12)
For the benchmark case where the mass of the DE mediator is on the order of the Hubble scale, $m_{\phi}\simeq H_{0}$, the present-day field value simplifies to $\bar{s}_{0}\simeq 1.45\beta^{1/2}$. This analytical estimate shows agreement with full numerical solutions of the cosmological evolution, as demonstrated in [Archidiacono:2022iuu]. Consequently, the effective mass correction for DM can be expressed directly in terms of the fundamental coupling strength $\beta$ by
$\Delta m_{\chi}^{2}\simeq 2.9\frac{H_{0}}{m_{\phi}}\beta^{1/2}m_{\chi}^{2}.$ (4.13)
@@
4.3 Applying superradiance constraints
@@
Figure 1: Normalised posterior probability density distributions for the DM mass $m_{\chi}$ and the coupling strength $\beta$, derived from observations of IRAS 09149-6206 (left) and M33 X-7 (right). The filled contours represent the results for our benchmark model with the DE mediator mass set to $m_{\phi}/H_{0}=1$, and the solid black lines indicate the corresponding 95% CRs at highest posterior density. For comparison, the dashed lines show the 95% CRs for models with alternative mediator masses of $m_{\phi}/H_{0}=2$ and $m_{\phi}/H_{0}=0.5$. Finally, the horizontal dashed line indicates the 95% C.L. cosmological upper limit on $\beta$ from Archidiacono et al. (2022) [Archidiacono:2022iuu].
To constrain the model parameters, we apply the statistical framework and methodology detailed in Section 3. We utilize the BH posterior samples redistributed by Hoof et al. (2024) [Hoof:2024quk], specifically focusing on two systems: the stellar-mass X-ray binary BH M33 X-7 [Liu:2008tk] and the SMBH IRAS 09149-6206 [GRAVITY:2020cli, Walton:2020vng, abuter2017first]. The analysis framework and associated code are based on the work presented in [Hoof:2024quk]  [^3] The code used for the analysis is publicly available at https://github.com/sebhoof/bhsr.. A crucial input for the superradiance constraint is the BH timescale, $\tau_{\mathrm{BH}}$, which represents the time available for the instability to develop. For accreting BHs, a conservative and physically motivated choice is the Salpeter time, $\tau_{\mathrm{Salp}}\approx 4.5\times 10^{7}$ yr, which represents the characteristic timescale for a BH to significantly increase its mass or spin up via Eddington-limited accretion (assuming a canonical 10% radiative efficiency, $\bar{\lambda}_{\mathrm{Edd}}=0.1$). For M33 X-7, we adopt its estimated age as the timescale ($\tau_{\mathrm{BH}}=\tau_{\mathrm{age}}=3\times 10^{6}\mathrm{yr}$ [Gou:2011nq]), while for the SMBH IRAS 09149-6206, we use the Eddington timescale ($\tau_{\mathrm{BH}}=\tau_{\mathrm{Edd}}=\tau_{\mathrm{Salp}}/\bar{\lambda}_{\mathrm{Edd}}=4.5\times 10^{8}\mathrm{yr}$ [Hoof:2024quk]), consistent with the Salpeter time.
@@
Our constraints can be directly compared with those from large-scale cosmological observations. The model presented here is identical to the Coupled Dark Energy model constrained by Archidiacono et al. (2022) [Archidiacono:2022iuu] using a combination of Cosmic Microwave Background data from Planck and Baryon Acoustic Oscillations data. Their analysis provides a $m_{\chi}$-independent upper bound on the coupling strength, $\beta\lesssim 0.00287$ at a 95% confidence level (C.L.). In Figure 1, we superimpose this cosmological constraint as a horizontal line.
The constraints presented are inherently two-dimensional, providing a joint exclusion on the $(m_{\chi},\beta)$ plane. To derive a specific limit on the coupling strength $\beta$ alone, prior knowledge of the DM mass $m_{\chi}$ is required. Fortunately, such information can be obtained from complementary cosmological and astrophysical observations. For example, studies of the Lyman-$\alpha$ forest provide robust lower bounds on the mass of ultralight scalar DM, as its quantum pressure suppresses the formation of small-scale structures. Current Lyman-$\alpha$ observations typically yield a lower bound $m_{\chi}\gtrsim 10^{-21}\mathrm{eV}$ for the scalar DM mass [Kobayashi:2017jcf]. By combining such independent constraints on $m_{\chi}$ with our superradiance results, it is possible to project the 2D exclusion region onto a one-dimensional constraint on the dark fifth force parameter $\beta$.
@@
5 Model II: DE superradiance induced by DM spike
@@
5.1 Non-minimally coupled quintessence model
@@
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[\frac{R}{16\pi G_{N}}-\frac{1}{2}g_{\mu\nu}\partial^{\mu}\Phi\partial^{\nu}\Phi-V(\Phi)\right]+S_{\mathrm{DM}}\left(\Psi;\Omega^{2}(\Phi)g_{\mu\nu}\right),$ (5.1)
@@
While models of this type are widely used in fifth-force [Uzan:2023dsk, Pitrou:2023swx, Hu:2025uqb] and cosmological studies [Wang:2025znm], the astrophysical behavior of the scalar field $\Phi$ in the strong gravity regime near BHs has also garnered significant attention within the context of scalar-tensor theories. It has been shown that surrounding matter can induce tachyonic instabilities leading to spontaneous scalarization or, relevantly for this work, trigger superradiant instabilities by providing an effective mass to the scalar field [cardosoMatterKerrBlack2013, Cardoso:2013fwa]. This phenomenon has been explored in various astrophysical environments, such as accretion disks [Lingetti:2022psy] and DM halos [Tanaka:2025bfl].
@@
5.2 DE superradiance induced by DM spike

### v2

@@ III.3 Statistical framework
To formalize this exclusion principle in a Bayesian context, we follow the statistical framework developed by Hoof et al. (2024) [67] but adapt it to our specific model and parameter space. The central idea is to compute the posterior probability for our model parameters $\boldsymbol{\alpha}$ given the observational data $D$ from a BH measurement.
@@
$p(\boldsymbol{\alpha}|D)$ $=p(\boldsymbol{\alpha})\int\!p(\boldsymbol{\beta}_{\mathrm{BH}}|D)\,p(\tilde{a}|M,\boldsymbol{\alpha})\,\mathrm{d}\boldsymbol{\beta}_{\mathrm{BH}}\,.$ (21)
@@
In practice, observational analyses provide a set of NN equally-weighted posterior samples, $\{\boldsymbol{\beta}_{\mathrm{BH}}^{i}=(M^{i},\tilde{a}^{i})\}_{i=1}^{N}$, for the BH parameters. The integral in Eq. (21) can therefore be accurately approximated via Monte Carlo integration as
@@
$p(\boldsymbol{\alpha}|D)\approx\frac{p(\boldsymbol{\alpha})}{N}\sum_{i=1}^{N}\Theta(\tilde{a}^{\text{crit}}(M^{i};\boldsymbol{\alpha})-\tilde{a}^{i})\,.$ (22)
@@
The sum effectively counts the fraction of posterior samples for the BH that are consistent with (i.e., not excluded by) the given set of model parameters $\boldsymbol{\alpha}$ [67]. Since the superradiance rate $\Gamma_{\mathrm{SR}}$, and thus $\tilde{a}^{\text{crit}}$, depends on the properties of the ultralight boson derived from the specific IDE model (such as its effective mass, which is a function of $\boldsymbol{\alpha}$), this statistical framework allows us to translate observational data of highly spinning BHs into exclusion regions in the parameter space $\boldsymbol{\alpha}$ of the IDE model under consideration.
IV Model I: DE as a dark fifth force mediator of the DM sector
@@
We now introduce the first of our specific field-theoretic models, building upon the general IDE framework discussed in Section II. In this scenario, the DE scalar field acts as a dark fifth force mediator of the DM sector, assuming both are ultralight scalar fields. This interaction alters the DM effective mass based on the background DE field, thereby modifying the superradiance potential of the DM particles themselves around spinning BHs and enabling constraints on the DE-DM coupling.
IV.1 Trilinear coupled quintessence model
@@
We begin by considering a coupled quintessence model where a DE scalar field, $\phi$, is coupled to an ultralight DM field  [^2] The assumption that DM is an ultralight scalar is motivated by the naturalness of the model [9]. Theoretical consistency requires that the parameters of the $\phi$-potential are at least as large as the radiative corrections induced by the interaction with DM. This leads to the condition $m_{\phi}^{2}\gtrsim\frac{\beta}{(4\pi)^{2}}\frac{m_{\chi}^{4}}{M_{\mathrm{Pl}}^{2}}\,.$ Imposing $m_{\phi}\lesssim H_{0}$ yields an upper bound on the DM mass, $m_{\chi}\lesssim\beta^{-1/4}\left(4\pi m_{\varphi}M_{\mathrm{Pl}}\right)^{1/2}\simeq 0.02\ \mathrm{eV}\left(\frac{0.01}{\beta}\right)^{1/4}\left(\frac{m_{\varphi}}{H_{0}}\right)^{1/2}\,.$ Therefore, a relativistic field-theoretical description of the dark fifth force mediator without fine-tuning requires the DM to be an ultralight boson. In this mass range, the DM candidate could be an axion or axion-like particle produced via the misalignment mechanism or other non-thermal processes., $\chi$, via a trilinear interaction term [53, 9, 26, 27]. The Lagrangian for this model is given by
$\mathcal{L}=-\frac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m_{\phi}^{2}\phi^{2}-\kappa\phi\chi^{2}\,$, (23)
@@
$\mathcal{L}=-\frac{1}{2}(\partial\chi)^{2}-\frac{1}{2}m_{\chi}^{2}(s)\chi^{2}-\frac{1}{2G_{s}}\left[(\partial s)^{2}+m_{\phi}^{2}s^{2}\right]\,.$ (25)
@@
Models of this type, where an ultralight scalar field (which may or may not be the DE field) mediates a long-range fifth force exclusively within the DM sector, have been extensively studied in the cosmological literature [125, 7, 53, 9, 26, 27]. These studies typically focus on constraining the model parameters using cosmological observations. Indeed, existing cosmological data, primarily from the Cosmic Microwave Background [9] and large-scale structure surveys [26, 27], already place stringent constraints on the dimensionless coupling strength, typically finding $\beta\lesssim\mathcal{O}(0.01)$. Our work complements these cosmological probes by exploring constraints from the distinct physical regime of BH superradiance.
IV.2 Effective mass correction
The coupling between the DE field $\phi$ and the DM field $\chi$ results in the mass of the DM particle becoming dependent on the background value of the DE field, $\bar{s}$. This gives rise to an effective mass for DM, which, for our model, is given by
$m_{\chi}^{2}(\bar{s})=m_{\chi}^{2}(1+2\bar{s}),$ (28)
where $m_{\chi}$ is the bare mass of the DM particle. [^3] It is crucial to emphasize that the DE mediator field $\phi$ (and consequently $s$) is characterized by an extraordinarily small bare mass, typically on the order of the present-day Hubble parameter ($m_{\phi}\sim H_{0}$). The corresponding Compton wavelength is on the cosmological scale ($\lambda_{\phi}\sim H_{0}^{-1}$). Because this macroscopic wavelength is astronomically larger than the localized scale of the BH, the field cannot resolve local spatial variations in the DM density. Therefore, the mediator field remains strictly homogeneous across the superradiant region, justifying our treatment of $\bar{s}$ as a spatially constant background value.
@@ IV.2 Effective mass correction
In principle, the value of the background field $\bar{s}$ evolves with cosmic time, governed by its equation of motion (see Eq (3.2) in [9]). A full analysis would require solving this equation over cosmic history, assuming a fixed initial value (for a scalar mass $m_{\phi}\sim H_{0}$, one finds $\bar{s}_{\mathrm{ini}}\sim\mathcal{O}(1)$). However, since the BH sample we selected formed very recently—roughly $\mathcal{O}(10-10^{2})\,\mathrm{Myr}$ ago, and the superradiance time scale is too short for significant change of $\bar{s}$, we can approximate $\bar{s}$ with its constant, present-day value $\bar{s}_{0}$ when applying superradiance constraints.
@@
$V(\bar{s}_{0})\approx\rho_{\mathrm{DE}}\approx 0.7\times 3H_{0}^{2}M_{\mathrm{Pl}}^{2}\,$, (29)
@@
$V(\bar{s})=\frac{1}{2}m_{\phi}^{2}\bar{\phi}^{2}=\frac{M_{\mathrm{Pl}}^{2}m_{\phi}^{2}\bar{s}^{2}}{\beta}.$ (30)
@@
$\bar{s}_{0}\approx 1.45\frac{H_{0}}{m_{\phi}}\beta^{1/2}.$ (31)
For the benchmark case where the mass of the DE mediator is on the order of the Hubble scale, $m_{\phi}\simeq H_{0}$, the present-day field value simplifies to $\bar{s}_{0}\simeq 1.45\beta^{1/2}$. This analytical estimate shows agreement with full numerical solutions of the cosmological evolution, as demonstrated in [9]. Consequently, the effective mass correction for DM can be expressed directly in terms of the fundamental coupling strength $\beta$ by
$\Delta m_{\chi}^{2}\simeq 2.9\frac{H_{0}}{m_{\phi}}\beta^{1/2}m_{\chi}^{2}.$ (32)
@@
IV.3 Applying superradiance constraints
@@
Figure 1: Normalised posterior probability density distributions for the DM mass $m_{\chi}$ and the coupling strength $\beta$, derived from observations of IRAS 09149-6206 (left) and M33 X-7 (right). The filled contours represent the results for our benchmark model with the DE mediator mass set to $m_{\phi}/H_{0}=1$, and the solid black lines indicate the corresponding 95% CRs at highest posterior density. For comparison, the dashed lines show the 95% CRs for models with alternative mediator masses of $m_{\phi}/H_{0}=2$ and $m_{\phi}/H_{0}=0.5$. Finally, the horizontal dashed line indicates the 95% C.L. cosmological upper limit on $\beta$ from Archidiacono et al. (2022) [9].
To constrain the model parameters, we apply the statistical framework and methodology detailed in Section III. We utilize the BH posterior samples redistributed by Hoof et al. (2024) [67], specifically focusing on two systems: the stellar-mass X-ray binary BH M33 X-7 [85] and the SMBH IRAS 09149-6206 [8, 120, 4]. The analysis framework and associated code are based on the work presented in [67]  [^4] The code used for the analysis is publicly available at https://github.com/sebhoof/bhsr.. A crucial input for the superradiance constraint is the BH timescale, $\tau_{\mathrm{BH}}$, which represents the time available for the instability to develop. For accreting BHs, a conservative and physically motivated choice is the Salpeter time, $\tau_{\mathrm{Salp}}\approx 4.5\times 10^{7}$ yr, which represents the characteristic timescale for a BH to significantly increase its mass or spin up via Eddington-limited accretion (assuming a canonical 10% radiative efficiency, $\bar{\lambda}_{\mathrm{Edd}}=0.1$). For M33 X-7, we adopt its estimated age as the timescale ($\tau_{\mathrm{BH}}=\tau_{\mathrm{age}}=3\times 10^{6}\mathrm{yr}$ [62]), while for the SMBH IRAS 09149-6206, we use the Eddington timescale ($\tau_{\mathrm{BH}}=\tau_{\mathrm{Edd}}=\tau_{\mathrm{Salp}}/\bar{\lambda}_{\mathrm{Edd}}=4.5\times 10^{8}\mathrm{yr}$ [67]), consistent with the Salpeter time.
@@
Our constraints can be directly compared with those from large-scale cosmological observations. The model presented here is identical to the Coupled Dark Energy model constrained by Archidiacono et al. (2022) [9] using a combination of Cosmic Microwave Background data from Planck and Baryon Acoustic Oscillations data. Their analysis provides a $m_{\chi}$-independent upper bound on the coupling strength, $\beta\lesssim 0.00287$ at a 95% confidence level (C.L.). In Figure 1, we superimpose this cosmological constraint as a horizontal line.
The constraints presented are inherently two-dimensional, providing a joint exclusion on the $(m_{\chi},\beta)$ plane. To derive a specific limit on the coupling strength $\beta$ alone, prior knowledge of the DM mass $m_{\chi}$ is required. Fortunately, such information can be obtained from complementary cosmological and astrophysical observations. For example, studies of the Lyman-$\alpha$ forest provide robust lower bounds on the mass of ultralight scalar DM, as its quantum pressure suppresses the formation of small-scale structures. Current Lyman-$\alpha$ observations typically yield a lower bound $m_{\chi}\gtrsim 10^{-21}\mathrm{eV}$ for the scalar DM mass [74]. By combining such independent constraints on $m_{\chi}$ with our superradiance results, it is possible to project the 2D exclusion region onto a one-dimensional constraint on the dark fifth force parameter $\beta$.
@@
V Model II: DE superradiance induced by DM spike
@@
V.1 Non-minimally coupled quintessence model
@@
$S$ $=\int\mathrm{d}^{4}x\sqrt{-g}\left[\frac{R}{16\pi G_{N}}-\frac{1}{2}g_{\mu\nu}\partial^{\mu}\Phi\partial^{\nu}\Phi-V(\Phi)\right]$
$+S_{\mathrm{DM}}\left(\Psi;\Omega^{2}(\Phi)g_{\mu\nu}\right)\,.$ (33)
@@
While models of this type are widely used in fifth-force [118, 96, 68] and cosmological studies [122], the astrophysical behavior of the scalar field $\Phi$ in the strong gravity regime near BHs has also garnered significant attention within the context of scalar-tensor theories. It has been shown that surrounding matter can induce tachyonic instabilities leading to spontaneous scalarization or, relevantly for this work, trigger superradiant instabilities by providing an effective mass to the scalar field [39, 38]. This phenomenon has been explored in various astrophysical environments, such as accretion disks [83] and DM halos [113].
@@
V.2 DE superradiance induced by DM spike